# QSO MUSEUM II: Search for extended Ly emission around 8 quasar pairs

### v1

@@ Acknowledgements.
Appendix A Example of a field before quasar PSF subtraction
@@
Appendix B Reanalyzed data of faint single quasar nebulae
In this work, we employ a conservative approach of using PNBs to detect extended Ly emission. Previously published samples of nebulae have often been optimally extracted within a 3D mask containing all pixels above a signal-to-noise ratio of 2 that are associated with the nebula. That method uses a less conservative detection criteria and it could therefore results in more extended detections. To facilitate comparison, we reanalyze the two most extreme nebulae published in Mackenzie et al. (2021) using 30 PNBs: ID4 is reported to host the faintest and smallest nebula (Fig. 15, left); ID6 is associated with the brightest and most extended Ly emission (Fig. 15, right). We find that our method is able to detect extended Ly emission as shown in Mackenzie et al. (2021), with the only exception of thin structures at low S/N and at the edges of the nebulae in their optimally extracted maps. For this reason, we find that our measurements for the extents (20 kpc for the small nebula and 134 kpc for the large nebula) are smaller than those reported in that work (60 kpc and 190 kpc). These differences are clearly due to the different extraction criteria for the nebulae but do not affect our results.

### v2

@@ Acknowledgements.
- Herwig et al. (2024) Herwig, E., Arrigoni Battaia, F., Bañados, E., & Farina, E. P. 2024, arXiv e-prints, arXiv:2410.00974
@@
Appendix A Additional material
In this appendix, we summarize important information about the MUSE observations and the quasar properties in Table 4 and show in Figs. 7 to 13 Ly maps and profiles for the remaining seven quasar pairs not shown in Fig. 3.
[…]
| 1.2 | 00:01:40.5984 | 07:09:47.8149 | 22.75 | -24.93 | 3.238 | ||||
| 2.2 | 00:18:08.0959 | 16:12:50.8420 | 22.13 | -25.42 | 3.126 | ||||
| 3.2 | 02:40:05.7477 | -00:39:14.0616 | 22.39 | -25.16 | 3.120 | ||||
[…]
| 6.2 | 10:21:16.9849 | 11:12:27.5716 | 20.28 | -28.01 | 3.815 | ||||
| 7.2 | 10:40:49.1050 | 01:00:33.0900 | 20.85 | -26.59 | 3.026 | ||||
| 8.2 | 11:35:02.5085 | -02:21:20.1432 | 22.06 | -25.36 | 3.009 |
Appendix B Example of a field before quasar PSF subtraction
@@
This becomes even more apparent when plotting the surface brightness profile of the data before subtraction, the constructed PSF and the data after subtraction in linearly spaced annuli as shown in Fig. 15 for PNBs of quasar pair 1. At the peak of the PSF, the unveiled extended emission is roughly 1.8 orders of magnitude dimmer than the quasar emission and it only starts dominating the emission in the PNB at a distance from the quasar of roughly 200 ckpc. We emphasize that during the analysis, we subtracted the empirical PSF channel-by-channel, and Fig. 15 therefore only functions as a simplified visualization of the subtraction.
Appendix C Reanalyzed data of faint single quasar nebulae
In this work, we employ a conservative approach of using PNBs to detect extended Ly emission. Previously published samples of nebulae have often been optimally extracted within a 3D mask containing all pixels above a signal-to-noise ratio of 2 that are associated with the nebula. That method uses a less conservative detection criteria, and it could therefore result in more extended detections. To facilitate comparison, we reanalyze the two most extreme nebulae published in Mackenzie et al. (2021) using 30 PNBs: ID4 is reported to host the faintest and smallest nebula (Fig. 16, left); ID6 is associated with the brightest and most extended Ly emission (Fig. 16, right). We find that our method is able to detect extended Ly emission as shown in Mackenzie et al. (2021), with the only exception of thin structures at low S/N and at the edges of the nebulae in their optimally extracted maps. For this reason, we find that our measurements for the extents (20 kpc for the small nebula and 134 kpc for the large nebula) are smaller than those reported in that work (60 kpc and 190 kpc). These differences are clearly due to the different extraction criteria for the nebulae but do not affect our results.